# appendixReferences

### v3

@@ 3.6 Computation
$\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{k}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{k}(\phi)},$ (25)
@@
$\max_{\phi\in\{0,1\}^{N}}v^{T}\phi+z_{1-\frac{\alpha}{2}}\sqrt{\phi^{T}Q\phi}$ (26)
where $\hat{\Phi}_{k}(\phi)=v^{T}\phi$ and $V_{k}(\phi)=\phi^{T}Q\phi$ for $v\in\mathbb{R}^{N}$ and $Q\in\mathbb{R}^{N\times N}$ given by
$v_{i}$ $=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{Y_{ki}\cdot P(Z_{i}=1)+(1-Y_{ki})\cdot P(Z_{i}=0)}$ $i\in[N]$ (27)
@@
in which $C_{ij}$ and $P_{ij}$ are given by the right hand sides of (23) and (24) respectively
@@
$\max_{\phi,u}$ $\ v^{T}\phi+z_{1-\frac{\alpha}{2}}\sqrt{\sum_{i=1}^{N}u_{i}}$
@@
It can be seen that $u_{i}=0$ if $\phi_{i}=0$, and $u_{i}=q_{i}^{T}\phi$ if $\phi_{i}=1$, enforcing that $\sum_{i}u_{i}=\phi^{T}Q\phi$. We solve this binary program using Gurobi through YALMIP (Lofberg, , 2004). For the instances that we considered, involving 1000 or more units, Gurobi was generally able to solve the problem to acceptable tolerances. If stopped early, Gurobi returns an upper bound, retaining the coverage properties of the actual solution.
@@ 4 Generalization to other estimands
We may generalize the approach of Section 3 to other estimands such as
  • The number of units affected by the treatment of others
  • The number of units affected by the treatment of others, excluding close neighbors
  • The number of units affected by the treatment of others when their direct treatment satisfies some condition
These estimands and their estimators will follow a general recipe.
  1. 1.
    First we introduce a vector $\phi^{*}\in\{0,1\}^{N}$ that identifies the units who do not meet the criteria of the estimand, and let $\Phi=\sum_{i=1}^{N}\phi_{i}^{*}$, so that the estimand is given by
    $\tau=N-\Phi$
  2. 2.
    We then introduce a pair of estimators for $\Phi$, which we denote by $\hat{\Phi}_{k}$ for $k=1,2$, or by $\hat{\Phi}_{k}^{{\textup{Haj}}}$ for their Hájek-weighted versions. A consistent lower bound for the estimand $\tau$ is given by
    $\hat{\tau}=N-\min\left\{\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)\,,\,\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)\right\}$ (29)
    where $\hat{\Phi}_{k}^{\textup{Haj}}(\phi)$ denotes $\hat{\Phi}_{k}^{\textup{Haj}}$, evaluated under the hypothesis that $\phi^{*}=\phi$.
  3. 3.
    If $\hat{\Phi}_{k}^{\textup{Haj}}$ for $k=1,2$ is consistent and asymptotically normal, with variance consistently estimated by $V_{k}$, then we can construct an asymptotic $(1-\alpha)$ level one-sided confidence interval for $\tau$, with lower boundary given by
    $N-\min\left\{\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{1}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{1}(\phi)}\,,\,\max_{\phi\in\{0,1\}^{N}}\hat{\Phi}_{2}^{\textup{Haj}}(\phi)+z_{1-\frac{\alpha}{2}}\sqrt{V_{2}(\phi)}\,\right\},$ (30)
    where $V_{k}(\phi)$ denotes $V_{k}$, evaluated under the hypothesis that $\phi^{*}=\phi$.
Following this recipe, we introduce the estimands $\tau^{\textup{indirect}}$, $\tau^{\textup{nonneighbors}}$, $\tau^{\textup{control}}$, and $\tau^{\textup{tr}}$, defining $(\phi^{*},\hat{\Phi}_{1},\hat{\Phi}_{2})$ for each one. After doing so, we present Hájek-normalized estimators and variance estimates in Sections 4.1 and 4.2.
The number of units affected by the treatment of others
Let $\phi_{i}^{*}\in\{0,1\}^{N}$ identify the units who are not affected by the treatment of others
$\phi_{i}^{*}=1\{\textup{if $Y_{i}$ is affected only by $X_{i}$, so that $Y_{i}=f_{i}(X_{i})$}\}$ (31)
@@
and let $\tau^{\textup{indirect}}=N-\Phi$, the number of units affected by the treatment of others. To estimate $\tau^{\textup{indirect}}$, we use $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ given by
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{W_{i}=Y_{ki}\}}{{P}(W_{i}=Y_{ki}|X_{i})}$
Long-range interference
@@
Let $\phi^{*}\in\{0,1\}^{N}$ identify the units who are affected only by their own treatment and that of their close neighbors,
$\phi_{i}^{*}=1\{\textup{if $Y_{i}$ is affected only by $X_{\eta_{i}}$, so that $Y_{i}=f_{i}(X_{\eta_{i}})$}\}$
@@
and let $\tau^{\textup{nonneighbors}}$ equal $N-\Phi$, the number of units affected by the treatment of others excluding close neighbors. To estimate $\tau^{\textup{nonneighbors}}$, we use $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ given by
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{W_{i}^{(2)}=Y_{ki}\}}{{P}(W_{i}^{(2)}=Y_{ki}|X_{\eta_{i}})}$
where $W_{i}^{(2)}$ is given by (3), and measures treatment of non-close neighbors.
@@
The number of units whose outcome under direct control (treatment) is affected by the treatment of others
@@
Let $\phi^{*}\in\{0,1\}^{N}$ identify the units whose outcome, when receiving control themselves, does not depend on the treatment of others
$\phi_{i}^{*}=1\{\textup{ if $Y_{i}$ is constant for all $X$ such that $X_{i}=0$}\}$
@@
in which case $\tau^{\textup{control}}=N-\Phi$ equals the number of units whose outcome under direct control is affected by the treatment of others. For example, if a vaccine grants immunity, then the disease status of a unit when receiving the placebo (but not necessarily when receiving the vaccine) might depend on the vaccination status of others.
To estimate $\tau^{\textup{control}}$, we consider $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ given by
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=0,W_{i}=Y_{ki}\}}{{P}(X_{i}=0,W_{i}=Y_{ki})}$
@@
We may similarly define $\tau^{\textup{tr}}$, the number of units whose outcome under direct treatment is affected by the treatment of others, by defining $\phi^{*}$ to be
$\phi_{i}^{*}$ $=1\{\textup{ if $Y_{i}$ is constant for all $X$ such that $X_{i}=1$}\}$
and $\hat{\Phi}_{1}$ and $\hat{\Phi}_{2}$ to be
$\hat{\Phi}_{k}$ $=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}=1,W_{i}=Y_{ki}\}}{{P}(X_{i}=1,W_{i}=Y_{ki})}$
@@ 4.1 General Functional Form
We give a general functional form for our estimands and estimators. In our estimands, $\phi^{*}\in\{0,1\}^{N}$ identifies the units whose outcome satisfies an exposure mapping condition for some subset of possible treatment assignments. We may write this as
$\phi_{i}^{*}=1\{\textup{ if $\exists\ \tilde{f}_{i}$ such that $f_{i}(X)=\tilde{f}_{i}(T_{i}(X))$ for all $X\in R_{i}$}\}$ (32)
where $T_{i}\equiv T_{i}(X)$ equals a vector of summary statistics describing unit $i$’s exposure to treatment. The set $R_{i}\subseteq\{0,1\}^{N}$ dictates when $T_{i}$ must be predictive of unit $i$’s outcome. For example, letting $T_{i}=X_{i}$ and $R_{i}=\{0,1\}^{N}$ induces $\phi^{*}$ given by (31).
The IPW estimators $\hat{\Phi}_{k}$ for $k=1,2$ are given by
$\hat{\Phi}_{k}=\sum_{i:\phi_{i}^{*}=1}\frac{1\{X_{i}\in R_{i},Z_{i}=Y_{ki}\}}{P(X_{i}\in R_{i},Z_{i}=Y_{ki}|T_{i})}$ (33)
for some choice of $Z_{i}\equiv Z_{i}(X)\in\{1,0,-1\}$ that maps unit $i$’s direct and indirect treatment to one of 3 groups (i.e., effective treatment, effective control, or not used).
Hájek-normalized versions of $\hat{\Phi}_{k}$ are given by
$\hat{\Phi}_{k}^{\textup{Haj}}=\sum_{i:\phi_{i}^{*}=1}\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{X\in R_{i},Z_{i}=Y_{ki}\}}{P(X\in R_{i},Z_{i}=Y_{ki}|T_{i})},$ (34)
@@
$\hat{N}_{a}$ $=\sum_{j=1}^{N}\frac{1\{X\in R_{i},Z_{i}=a\}}{P(X\in R_{i},Z_{i}=a|T_{i})},\qquad a\in\{0,1\}.$ (35)
As will be shown in Theorem 4, using the Hájek-weighted estimates to construct $\hat{\tau}$ by (29) will result in a Hájek-weighted comparison of outcomes, given by
$\hat{\tau}=\left|\frac{N}{\hat{N}_{1}}\cdot\frac{1\{X\in R_{i},Z_{i}=1\}}{P(X\in R_{i},Z_{i}=1|T_{i})}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\cdot\frac{1\{X\in R_{i},Z_{i}=0\}}{P(X\in R_{i},Z_{i}=0|T_{i})}Y_{i}\right|,$ (36)
Table 1 gives values for $\{T_{i},R_{i},Z_{i}\}_{i=1}^{N}$ corresponding to the estimands considered in this paper, while Table 2 gives expressions for the contrasts given by (36).
@@
$\tau^{\textup{basic}}$ (w/network) 1 $\{0,1\}^{N}$ $1\{X_{i}=W_{i}=1\}-1\{X_{i}\neq W_{i}\}$
@@
Table 1: Exposure mappings, subsets, and contrast mappings for various estimands. Note that letting $T_{i}=1$ for all $X$ enforces $Y_{i}$ to equal a constant for all $X\in R_{i}$, should $\phi_{i}^{*}=1$.
Estimand Hájek-normalized $\hat{\tau}$, expressed as contrast given by (36)
$\tau^{\textup{basic}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{X_{i}}{P(X_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1-X_{i}}{P(X_{i}=0)}Y_{i}\right|}$
$\tau^{\textup{basic}}$ (w/network) ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=W_{i}=1\}}{P(X_{i}=W_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=W_{i}=0\}}{P(X_{i}=W_{i}=0)}Y_{i}\right|}$
$\tau^{\textup{indirect}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{W_{i}=1\}}{P(W_{i}=1|X_{i})}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{W_{i}=0\}}{P(W_{i}=0|X_{i})}Y_{i}\right|}$
$\tau^{\textup{nonneighbors}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{W_{i}^{(2)}=1\}}{P(W_{i}^{(2)}=1|X_{\eta_{i}})}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{W_{i}^{(2)}=0\}}{P(W_{i}^{(2)}=0|X_{\eta_{i}})}Y_{i}\right|}$
$\tau^{\textup{control}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=0,W_{i}=1\}}{P(X_{i}=0,W_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=0,W_{i}=0\}}{P(X_{i}=0,W_{i}=0)}Y_{i}\right|}$
$\tau^{\textup{tr}}$ ${\left|\sum_{i=1}^{N}\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=1,W_{i}=1\}}{P(X_{i}=1,W_{i}=1)}Y_{i}-\sum_{i=1}^{N}\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=1,W_{i}=0\}}{P(X_{i}=1,W_{i}=0)}Y_{i}\right|}$
@@
4.2 Variance Estimation for General Case

### v4

@@ 3.7 Computation
$\min_{\phi\in\{0,1\}^{N}}\hat{\tau}_{k}^{\textup{Haj}}(\phi)-z_{1-\frac{\alpha}{2}}\sqrt{\hat{V}_{k}(\phi)},$ (22)
@@
$\min_{\phi\in\{0,1\}^{N}}N-b^{T}\phi-z_{1-\frac{\alpha}{2}}\sqrt{\phi^{T}Q\phi}$ (23)
where $\hat{\tau}_{k}^{\textup{Haj}}(\phi)=N-b^{T}\phi$ and $\hat{V}_{k}(\phi)=\phi^{T}Q\phi$ for $b\in\mathbb{R}^{N}$ and $Q\in\mathbb{R}^{N\times N}$ given by
$b_{i}$ $=\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{Z_{i}=Y_{ki}\}}{Y_{ki}\cdot P(Z_{i}=1)+(1-Y_{ki})\cdot P(Z_{i}=0)}$ $i\in[N]$ (24)
@@
in which $C_{ij}$ and $P_{ij}$ appearing in (25) are given by the right hand sides of (20) and (21) respectively
@@
$\max_{\phi,u}$ $\ b^{T}\phi+z_{1-\frac{\alpha}{2}}\sqrt{\sum_{i=1}^{N}u_{i}}$
@@
It can be seen that $u_{i}=0$ if $\phi_{i}=0$, and $u_{i}=q_{i}^{T}\phi$ if $\phi_{i}=1$, enforcing that $\sum_{i}u_{i}=\phi^{T}Q\phi$. We solve this binary program using Gurobi, a commercial optimization solver. For the instances that we considered, involving 1000 or more units, Gurobi was generally able to solve the problem to acceptable tolerances. If stopped early, Gurobi returns an upper bound, retaining the coverage properties of the actual solution.
@@ 4 Generalization to other estimands
We now generalize the approach of Section 3 to cover estimands that more directly characterize interference, such as those described in the introduction. In Section 4.1 we introduce the estimands $\tau^{\textup{indirect}}$, $\tau^{\textup{nonneighbors}}$, $\tau^{\textup{control}}$, and $\tau^{\textup{tr}}$, defining $\mathcal{I}$} and Horvitz-Thompson estimators for each one. In Section 4.2 we then give a general formulation that covers each of these examples, and define the Hájek-normalized estimators $\hat{\tau}_{1}^{\textup{Haj}}$ and $\hat{\tau}_{2}^{\textup{Haj}}$ and the Hájek-normalized contrast $\Delta^{\textup{Haj}}$. Section 4.3 then presents variance estimators and inference.
4.1 Other Estimands
The number of units affected by the treatment of others ($\tau^{\textup{indirect}}$)
Let ${\mathcal{I}}\subseteq[N]$ identify the units who are not affected by the treatment of others
[…]
@@ 4.1 Other Estimands
and let $\tau^{\textup{indirect}}=N-|{\mathcal{I}}|$, the number of units affected by the treatment of others.
To estimate $\tau^{\textup{indirect}}$, we may use estimators $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ given by
$\hat{\tau}_{k}$ $=N-\sum_{i\in{\mathcal{I}}}\frac{1\{W_{i}=Y_{ki}\}}{{P}(W_{i}=Y_{ki}|X_{i})}.$
This estimator differs from (13) in that we condition on $X_{i}$ in the probability weight $P(W_{i}=Y_{ki}|X_{i})$. This is done because the outcome $Y_{ki}$ is no longer constant for $i\in{\mathcal{I}}$, but instead is a function of $X_{i}$. As a result, the probability $P(W_{i}=Y_{ki})$ is unknown if $f_{i}$ is unknown, even if $i\in{\mathcal{I}}$. On the other hand, the conditional probability $P(W_{i}=Y_{ki}|X_{i})$ may be evaluated without knowledge of $f_{i}$, according to
$P(W_{i}=Y_{ki}|X_{i})=Y_{ki}\cdot P(W_{i}=1|X_{i})+(1-Y_{ki})\cdot P(W_{i}=0|X_{i}),\qquad i\in{\mathcal{I}}$
as will be shown by Proposition 4 in Section 4.2.
Long-range interference ($\tau^{\textup{nonneighbors}}$)
@@
Let ${\mathcal{I}}\subseteq[N]$ identify the units who are affected only by their own treatment and that of their close associates,
[…]
@@
and let $\tau^{\textup{nonneighbors}}$ equal $N-|{\mathcal{I}}|$, the number of units affected by the treatment of others excluding close associates. To estimate $\tau^{\textup{nonneighbors}}$, we may use $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ given by
$\hat{\tau}_{k}$ $=N-\sum_{i\in{\mathcal{I}}}\frac{1\{W_{i}^{(2)}=Y_{ki}\}}{{P}(W_{i}^{(2)}=Y_{ki}|X_{{\mathcal{N}}_{i}})}$
where $W_{i}^{(2)}$ is given by (3), and measures treatment of non-close associates.
@@
The number of units whose outcome under direct control (treatment) is affected by the treatment of others ($\tau^{\textup{control}}$, $\tau^{\textup{tr}}$)
@@
Let ${\mathcal{I}}\subseteq[N]$ identify the units whose outcome, when receiving control themselves, does not depend on the treatment of others
[…]
@@
in which case $\tau^{\textup{control}}=N-|{\mathcal{I}}|$ equals the number of units whose outcome under direct control is affected by the treatment of others. For example, if a vaccine grants immunity, then the disease status of a unit when receiving the placebo (but not necessarily when receiving the vaccine) might depend on the vaccination status of others.
To estimate $\tau^{\textup{control}}$, we consider $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ given by
$\hat{\tau}_{k}$ $=N-\sum_{i\in{\mathcal{I}}}\frac{1\{X_{i}=0,W_{i}=Y_{ki}\}}{{P}(X_{i}=0,W_{i}=Y_{ki})}$
@@
We may similarly define $\tau^{\textup{tr}}$, the number of units whose outcome under direct treatment is affected by the treatment of others, by defining $\mathcal{I}$} to be
[…]
and $\hat{\tau}_{1}$ and $\hat{\tau}_{2}$ to be
$\hat{\tau}_{k}$ $=N-\sum_{i\in{\mathcal{I}}}\frac{1\{X_{i}=1,W_{i}=Y_{ki}\}}{{P}(X_{i}=1,W_{i}=Y_{ki})}$
@@ 4.2 General Functional Form
We give a general functional form for our estimands and estimators. We may write our estimands as
$\tau=N-|{\mathcal{I}}|,$ (27)
where ${\mathcal{I}}\subseteq[N]$ contains the units whose outcome mapping satisfies an interference specification, which may be written as
${\mathcal{I}}=\{i:\textup{ $\exists\ \tilde{f}_{i}$ such that $f_{i}(X)=\tilde{f}_{i}(T_{i}(X))$ for all $X\in R_{i}$}\}$ (28)
where $T_{i}\equiv T_{i}(X)$ equals a vector of summary statistics describing unit $i$’s exposure to treatment, and $R_{i}$ is a subset of $\{0,1\}^{N}$. For units in $\mathcal{I}$}, the value of $T_{i}$ dictates unit $i$’s outcome whenever $X\in R_{i}$. Both $T_{i}$ and $R_{i}$ are chosen by the researcher to define the specification of interest. For example, letting $T_{i}=X_{i}$ and $R_{i}=\{0,1\}^{N}$ induces $\mathcal{I}$} given by (26).
In addition to $T_{i}$ and $R_{i}$, the researcher also chooses $Z_{i}\equiv Z_{i}(X)\in\{1,0,-1\}$, which maps unit $i$’s treatment exposure to one of 3 groups (i.e., effect treatment, effective control, or not used). Table 1 gives values for $\{T_{i},R_{i},Z_{i}\}_{i=1}^{N}$ corresponding to the estimands considered in this paper.
To simplify notation, let $\bar{Z}_{i}$ for $i\in[N]$ equal $Z_{i}$ if $X\in R_{i}$, and equal -1 otherwise,
The Horvitz-Thompson estimators $\hat{\tau}_{k}$ for $k=1,2$ are given by
$\hat{\tau}_{k}=N-\sum_{i\in{\mathcal{I}}}\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})}$ (30)
and Hájek-normalized versions of $\hat{\tau}_{k}$ are given by
$\hat{\tau}_{k}^{\textup{Haj}}=N-\sum_{i\in{\mathcal{I}}}\frac{N}{\hat{N}_{Y_{ki}}}\cdot\frac{1\{\bar{Z}_{i}=Y_{ki}\}}{P(\bar{Z}_{i}=Y_{ki}|T_{i})},$ (31)
@@
$\hat{N}_{a}$ $=\sum_{i=1}^{N}\frac{1\{\bar{Z}_{i}=a\}}{P(\bar{Z}_{i}=a|T_{i})},\qquad a\in\{0,1\}.$ (32)
As will be shown in Theorem 1, the Hájek-weighted estimates will be lower bounded by the absolute value of a Hájek-weighted comparison of outcomes given by
$\Delta^{{\textup{Haj}}}=\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}-\frac{N}{\hat{N}_{0}}\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\right)Y_{i}$ (33)
For the estimands considered in this paper, Table 2 gives expressions for $\Delta^{{\textup{Haj}}}$.
@@
$\tau^{\textup{basic}}$ using (9) 1 $\{0,1\}^{N}$ $1\{X_{i}=W_{i}=1\}-1\{X_{i}\neq W_{i}\}$
@@
Table 1: Exposure mappings, subsets, and contrast mappings for various estimands. Note that letting $T_{i}=1$ for all $X$ enforces $Y_{i}$ to equal a constant for all $X\in R_{i}$, if $i\in{\mathcal{I}}$.
Estimand Hájek-normalized contrast $\Delta^{\textup{Haj}}$
$\tau^{\textup{basic}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{X_{i}}{P(X_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{1-X_{i}}{P(X_{i}=0)}\right)Y_{i}}$
$\tau^{\textup{basic}}$ (w/network) ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=W_{i}=1\}}{P(X_{i}=W_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=W_{i}=0\}}{P(X_{i}=W_{i}=0)}\right)Y_{i}}$
$\tau^{\textup{indirect}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{W_{i}=1\}}{P(W_{i}=1|X_{i})}-\frac{N}{\hat{N}_{0}}\frac{1\{W_{i}=0\}}{P(W_{i}=0|X_{i})}\right)Y_{i}}$
$\tau^{\textup{nonneighbors}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{W_{i}^{(2)}=1\}}{P(W_{i}^{(2)}=1|X_{{\mathcal{N}}_{i}})}-\frac{N}{\hat{N}_{0}}\frac{1\{W_{i}^{(2)}=0\}}{P(W_{i}^{(2)}=0|X_{{\mathcal{N}}_{i}})}\right)Y_{i}}$
$\tau^{\textup{control}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=0,W_{i}=1\}}{P(X_{i}=0,W_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=0,W_{i}=0\}}{P(X_{i}=0,W_{i}=0)}\right)Y_{i}}$
$\tau^{\textup{tr}}$ ${\sum_{i=1}^{N}\left(\frac{N}{\hat{N}_{1}}\frac{1\{X_{i}=1,W_{i}=1\}}{P(X_{i}=1,W_{i}=1)}-\frac{N}{\hat{N}_{0}}\frac{1\{X_{i}=1,W_{i}=0\}}{P(X_{i}=1,W_{i}=0)}\right)Y_{i}}$
@@
Analogous to the discussion of Section 3, Proposition 3 gives intuition for when our lower bounds may be tight, and suggests guidance for the choice of $\bar{Z}$. It states that if for all $i\notin{\mathcal{I}}$ the contrast variable $\bar{Z}_{i}$ is a perfect predictor of the outcome $Y_{i}$ whenever $\bar{Z}_{i}\in\{0,1\}$, then an IPW weighted contrast using $\bar{Z}_{i}$ will be unbiased for the estimand.
Proposition 3.
Let $\Delta^{\bar{Z}}$ equal
$\Delta^{\bar{Z}}=\sum_{i=1}^{N}\left(\frac{1\{\bar{Z}_{i}=1\}}{P(\bar{Z}_{i}=1|T_{i})}-\frac{1\{\bar{Z}_{i}=0\}}{P(\bar{Z}_{i}=0|T_{i})}\right)Y_{i}$
and for all $i\notin{\mathcal{I}}$ let $Y_{i}=\bar{Z}_{i}$ whenever $\bar{Z}_{i}\neq-1$. Then $\mathbb{E}[\Delta^{\bar{Z}}]=N-|{\mathcal{I}}|.$
4.3 Variance Estimation and Inference for General Case